# Auto-Calibration and Biconvex Compressive Sensing with Applications to Parallel MRI

### v1

@@ 2.3 Our contributions
We propose a new method that combines convex optimization and compressive sensing to solve the auto-calibration problem in the form of (2) by exploiting the block sparsity structure within the lifted signals among all parallel sensors and solving a convenient mixed-norm optimization problem. The main contribution is to provide a robust and stable theoretical guarantee of recovery. Additionally, existing theoretical development in calibrationless compressive sensing methods often lacks assumptions aligned with those suitable for the pMRI application. We make an effort to incorporate appropriate assumptions throughout our analysis.
@@ 3 Problem setup
where $\textbf{y}_{i}\in\mathbb{C}^{L\times 1}$ are the measurements, $\textbf{F}_{\Omega}\in\mathbb{C}^{L\times N}$ is the partial Tourier matrix of the N-point DFT matrix with $L$ rows chosen from the subset $\Omega\in[N]$ , $\mathbf{D}_{i}\in\mathbb{C}^{N\times N}$ is the diagonal matrices of calibration parameters, $\textbf{x}\in\mathbb{C}^{N\times 1}$ the underlying signal, $\textbf{n}_{i}\in\mathbb{C}^{N\times 1}$ being additive noise.
Our goal is to recover the common signal of interest $\mathbf{x}$ from the measurements $\mathbf{y}_{i}$ where the calibration matrices $\mathbf{D}_{i}$ are unknown. As discussed in the introduction, it is in general not possible to recover x (and $\textbf{D}_{i}$) from the measurements since the number of unknowns is larger than the number of measurements, regardless of how many independent experiments are taken (reflected in the number $C$).
Hence, we impose two assumptions on both the signal $\mathbf{x}$ and the calibration parameters $\textbf{D}_{i}$. In particular,
@@
    x has a sparse representation: there is a known sparsifying transformation $\mathbf{\Psi}\in\mathbb{C}^{N\times N}$ such that $\mathbf{z}=\mathbf{\Psi}\mathbf{x}$ is n-sparse. The location of the $n$ active coefficient of z is unknown.
@@
    The calibration matrices belong to a common and known subspace: there is a known tall matrix $\textbf{B}\in\mathbb{C}^{N\times k}$ with $k<N$ such that $\mathbf{D}_{i}=\text{diag}(\mathbf{B}\textbf{h}_{i}),\text{for all }i=1\cdots C$. Here $\mathbf{h_{i}}\in\mathbb{C}^{k}$ are unknown.
@@
Next, we apply the idea of ‘Lifting’ to transform the bi-linear constraint in both $\textbf{h}_{i}$ and z to a linear constraint with respect to their outer product $\textbf{h}_{i}\textbf{z}^{T}$
@@ 3.1 Lifting
Using the lifting approach [11, 4, 36], we can transform the biconvex problem into the problem of recovering a sparse matrix from linear measurements.
In particular, denote $\textbf{B}=\begin{bmatrix}\mathbf{b}_{1}^{*}\\ \vdots\\ \mathbf{b}_{N}^{*}\end{bmatrix}=\begin{bmatrix}\mathbf{c}_{1}&\cdots&\mathbf{c}_{k}\end{bmatrix}\in\mathbb{C}^{N\times k}$, $\mathbb{\Psi}^{*}=\begin{bmatrix}\mathbf{q}_{1}^{T}\\ \vdots\\ \mathbf{q}_{N}^{T}\end{bmatrix}=\begin{bmatrix}\boldsymbol{\psi}_{1}&\cdots&\boldsymbol{\psi}_{N}\end{bmatrix}\in\mathbb{C}^{N\times N}$.
@@
$\textbf{y}_{i}=\textbf{F}_{\Omega}\begin{bmatrix}\textbf{b}_{1}^{*}\textbf{h}_{i}\textbf{z}^{T}\textbf{q}_{1}\\ \textbf{b}_{2}^{*}\textbf{h}_{i}\textbf{z}^{T}\textbf{q}_{2}\\ \vdots\\ \textbf{b}_{N}^{*}\textbf{h}_{i}\textbf{z}^{T}\textbf{q}_{N}\end{bmatrix}=\mathbb{F_{\Omega}}\underbrace{\begin{bmatrix}(\mathbf{\bar{q}}_{1}\bigotimes\mathbf{b}_{1})^{*}\\ \vdots\\ \mathbf{\bar{q}}_{N}\bigotimes\mathbf{b}_{N})^{*}\end{bmatrix}}_{\mathbf{\Phi}}\text{vec}(\textbf{h}_{i}\textbf{z}^{T}):=\mathbf{F}_{\Omega}\mathbf{\Phi}\text{vec}(\textbf{h}_{i}\textbf{z}^{T}),i=1\cdots C,$ (13)
where $\boldsymbol{\Phi}=\begin{bmatrix}(\mathbf{\bar{q}}_{1}\bigotimes\mathbf{b}_{1})^{*}\\ \vdots\\ \mathbf{\bar{q}}_{N}\bigotimes\mathbf{b}_{N})^{*}\end{bmatrix}\in\mathbb{C}^{N\times kN}$.
Define $\textbf{Y}=\begin{bmatrix}\textbf{y}_{1}|\textbf{y}_{2}|\cdots|\textbf{y}_{C}\end{bmatrix}$, $\textbf{H}=\begin{bmatrix}\textbf{h}_{1}|\textbf{h}_{2}|\cdots|\textbf{h}_{C}\end{bmatrix}$, $\textbf{X}=\begin{bmatrix}\textbf{X}_{1}|\textbf{X}_{2}|\cdots|\textbf{X}_{C}\end{bmatrix}=\begin{bmatrix}\text{vec}(\textbf{h}_{1}\textbf{z}^{T})|\text{vec}(\textbf{h}_{2}\textbf{z}^{T})|\cdots|\text{vec}(\textbf{h}_{C}\textbf{z}^{T})\end{bmatrix}=\textbf{z}\otimes\textbf{H}\in\mathbf{C}^{kN\times C}$, we can rewrite (13) in matrix form as follows:
@@
where $\textbf{Y}\in\mathbb{C}^{L\times C}$, $\mathbf{F}_{\Omega}\in\mathbb{C}^{L\times N}$, $\mathbf{\Phi}\in\mathbb{C}^{N\times kN}$, $\textbf{X}\in\mathbb{C}^{KN\times 1}$. We refer to $\mathbf{A}=\mathbf{F}_{\Omega}\mathbf{\Phi}$ as the forward matrix. Now the problem is to find the matrix $\mathbf{X}$ satisfying the linear constraint (14).
@@ 3.2 Block sparsity of $\mathbf{X}$
Moreover, X is also low rank. Consider a reshaped matrix $\mathbf{\tilde{X}}=\begin{bmatrix}\textbf{h}_{1}\textbf{z}^{T}&\cdots&\textbf{h}_{c}\textbf{z}^{T}\end{bmatrix}^{T}\in\mathbb{C}^{kC\times N}$, which has rank 1. Methods that exploit the low-rankness are introduced in [5][52][27] and include the ENLIVE formulation (8). Naturally, one could minimize over a linear combination $\|\textbf{X}\|_{1,2}+\lambda\|\textbf{X}\|_{*},\lambda>0$ in an attempt to simultaneously capture the low-rankness and sparse characteristics of X. However, as demonstrated in reference [43], this approach is no more effective than using a single norm in terms of the number of measurements required. Theoretically, only recovering for a single coil case by minimizing $*\|\cdot\|_{*}$ requires $L=O(N+k)$. Also $l_{1,2}$ minimization is more computationally efficient than the nuclear norm minimization problem. Hence, we choose to only minimize with respect to $\|\cdot\|_{1,2}$, which (as we will prove) may already yield a solution with sparse structure and small nuclear norm.
@@ 3.3 Optimization problem
which is the sum of the Frobenious norm of the $k$ by $C$ blocks of $\mathbf{X}$.
Since it is well known that the $\ell_{1,2}$-norm promotes the block sparse structure in $\mathbf{X}$ [17][18][20], we solve the following convex program
@@
If the measurement are noisy, i.e., $\textbf{Y}=\textbf{F}_{\Omega}\mathbf{\Phi}\textbf{X}+\mathbf{N}$ with $\|\mathbf{N}\|_{F}\leq\sigma$, we solve the following convex program
$\min\|\mathbf{X}\|_{1,2}\quad\text{subject to}\|\textbf{Y}-\textbf{F}_{\Omega}\mathbf{\Phi}\textbf{X}\|_{F}\leq\sigma.$ (17)
@@
Once we get a solution $\hat{\textbf{X}}$ to the $\ell_{1,2}$ minimization problem (16)(17), we calculate the average of the best rank-one estimations of its columns to retrieve the underlying signal. Recall the columns of $\hat{\textbf{X}}=\begin{bmatrix}\hat{\textbf{x}}_{1}&\cdots\hat{\textbf{x}}_{C}\end{bmatrix}$ and the truth $\textbf{X}^{0}=\begin{bmatrix}\textbf{X}^{0}_{1}&\cdots\textbf{X}^{0}_{C}\end{bmatrix}=\begin{bmatrix}\text{vec}(\textbf{h}_{1}\textbf{z}^{T})|\text{vec}(\textbf{h}_{2}\textbf{z}^{T})|\cdots|\text{vec}(\textbf{h}_{C}\textbf{z}^{T})\end{bmatrix}$. We first reshape the columns $\hat{\textbf{x}}_{i}$ into a matrix $\hat{\textbf{X}}_{i}\in\mathbb{C}^{k\times N}$ in the column-wise order, i.e., $\text{vec}(\hat{\textbf{X}}_{i})=\hat{\textbf{x}}_{i}$. Consider the best rank-one estimation of $\hat{\textbf{X}}_{i}$, $\sigma_{i}\hat{\textbf{u}}_{i}\hat{\textbf{v}}_{i}^{T}$, where $\sigma_{i}$, $\hat{\textbf{u}}_{i}$ and $\hat{\textbf{v}}_{i}$ are its leading singular value, left and right singular vectors respectively. Let $\hat{\textbf{h}}_{i}=\sqrt{\sigma_{i}}\hat{\textbf{u}}_{i}$ and $\hat{\textbf{z}}_{i}=\sqrt{\sigma_{i}}\hat{\textbf{v}}_{i}$. We use the following lemma from [36] to ensure that $\hat{\textbf{h}}_{i}$ and $\hat{\textbf{z}}_{i}$ are close to $\textbf{h}_{i}$ and z.
@@
For any matrix X and $\textbf{X}_{0}=\textbf{u}_{0}\textbf{v}_{0}^{T}$, let $\hat{\sigma}\hat{\textbf{u}}\hat{\textbf{v}}^{T}$ be the best rank-one Frobenius norm approximation of X. If $\|\textbf{X}-\textbf{X}_{0}\|_{F}=\epsilon$, then there exists a scalar $\alpha_{0}$ and a constant $C_{0}$ such that,
$\|\textbf{u}_{0}-\alpha_{0}\hat{\textbf{u}}\|\leq C_{0}\min(\epsilon/\|\textbf{u}_{0}\|,\|\textbf{u}_{0}\|),\|\textbf{v}_{0}-\alpha_{0}^{-1}\hat{\textbf{v}}\|\leq C_{0}\min(\epsilon/\|\textbf{v}_{0}\|,\|\textbf{v}_{0}\|).$
Notice if $\textbf{X}_{0}=\textbf{u}_{0}\textbf{v}_{0}^{T}$, then $\alpha_{0}\textbf{u}_{0}$ and $1/\alpha_{0}\textbf{v}_{0}$ is also a pair of solutions for any $\alpha_{0}\neq 0$. However, the scalar ambiguity is not important for many applications, including MRI. We set our estimation of z as an average $\hat{\textbf{z}}=\sum\limits_{i=1\cdots C}\hat{\textbf{z}}_{i}/C$. The reconstructed signal is $\hat{\textbf{x}}=\boldsymbol{\Psi}^{*}\hat{\textbf{z}}$
@@
A comprehensive review of the algorithms to solve problems of the form (18) is not a focus of this paper; we refer to a few common approaches such as the forward-backward splitting (FBS) [14], ADMM [55] and FISTA [8] which have found numerous applications in the context of pMRI reconstruction.
@@ 4 Main Result
Given any integer $k$ and $N$, we define $T_{j}\subset[kN]$ by $T_{j}=\{(j-1)k+1,\cdots,jk\},j\in[N]$. Given an index set $S=\{s_{1},s_{2},\cdots\}\in[N]$, define $S_{j}=T_{s_{j}}$ for $j\in|S|$ and $T:=\cup_{j\in|S|}S_{j}$. With slight abuse of notation, we will not distinguish $\mathcal{P}_{S}$, $\mathcal{P}_{T}$ and $\mathcal{P}_{T\times[C]}$ when it is clear form the context. For example, for $\textbf{x}\in C^{N}$, $\textbf{x}\in C^{kN}$ and $\textbf{X}\in C^{kN\times C}$, $\mathcal{P}_{S}$ means the projection onto $S$, $T$ and $T\times[C]$ respectively. We will also sometimes write the projections $\mathcal{P}_{S}$ as subscripts, for example, $\textbf{x}_{S}$ or $\textbf{X}_{S}$, when it is convenient.
@@
It will come handy in the proofs to decompose a linear maps to any given subset of its columns. For a matrix $\textbf{A}\in\mathbb{C}^{L\times kN}$ and a given set $S$, we define the projections onto its blocks of columns induced by $S$ as

### v2

@@ 2.3 Our contributions
While the aforementioned auto-calibration methods in Section 2.2.1 undoubtedly offer utility and applicability by giving successful numerical results, the majority of them do not provide rigorous analysis or theoretical guarantees of recovery. Note that even in the known calibration case (14), the incoherence of the sensing matrix has not been properly analyzed for the popular $\ell_{1}$ minimization scheme, particularly for measurement ensembles that are applicable to pMRI. In this work, we take a step forward to provide analysis for the more complicated auto-calibration problem in (1) and to provide sufficient conditions on $L$ and $C$ for stable and robust recovery of an approximate transform sparse signal x.
We propose to solve the auto-calibration problem in (2) using convex optimization by leveraging the block sparsity structure across all channels simultaneously. Our approach efficiently solves a mixed norm minimization problem and provides robust and stable recovery with theoretical guarantees. Interestingly, our theoretical findings align with previous numerical observations from past literature, providing valuable explanatory insights. To provide additional evidence, we present numerical results using the proposed method to solve the auto-calibrated pMRI. We include simulations using the analytical phantom and simulated coil sensitivities, as well as reconstructions from real pMRI measurements.
@@ 3 Problem setup
where $\textbf{y}_{i}\in\mathbb{C}^{L\times 1}$ are the measurements, $\textbf{F}_{\Omega}\in\mathbb{C}^{L\times N}$ is the partial Fourier matrix with its rows chosen from the N-point 2-dimensional discrete Fourier transformation (DFT) matrix which belongs to a subset $\Omega\subset[N]$ , $\mathbf{D}_{i}=\text{diag}(\textbf{s}_{i})\in\mathbb{C}^{N\times N}$ is the diagonal matrix of calibration parameters corresponding to the $i$-th channel/coil, $\textbf{x}\in\mathbb{C}^{N\times 1}$ is the signal of interest and $\textbf{n}_{i}\in\mathbb{C}^{N\times 1}$ is additive noise.
Our goal is to recover the common signal of interest $\mathbf{x}$ from the measurements $\mathbf{y}_{i}$ across all channels, while the calibration matrices $\mathbf{D}_{i}$ are unknown. As discussed in the previous Sections, such a task is impossible unless additional assumptions are made on the signal and/or the calibration parameters.
Hence, we adopt the following two assumptions on both the signal $\mathbf{x}$ and the calibration parameters $\textbf{D}_{i}$, as justified in the previous discussions. In particular,
@@
    x has a sparse representation: there is a known sparsifying transformation $\mathbf{\Psi}\in\mathbb{C}^{N\times N}$ such that $\mathbf{z}=\mathbf{\Psi}\mathbf{x}$ is n-sparse. The location of the $n$ active coefficient of z is unknown;
@@
    The calibration matrices belong to a common and known subspace: there is a known tall matrix $\textbf{B}\in\mathbb{C}^{N\times k}$ with $k<N$ such that $\mathbf{D}_{i}=\text{diag}(\mathbf{B}\textbf{h}_{i}),\text{for all }i=1\cdots C$. Here, $\mathbf{h_{i}}\in\mathbb{C}^{k}$ are unknown.
@@
Next, we apply the idea of ‘lifting’ to transform the bi-linear constraint in both $\textbf{h}_{i}$ and z to a linear constraint with respect to their outer product $\textbf{h}_{i}\textbf{z}^{T}$.
@@ 3.1 Lifting
Using the lifting approach [Candès_Eldar_Strohmer_Voroninski_2013, Ahmed_Recht_Romberg_2014, ling_strohmer_2015], we can transform the biconvex constraint (22) into the problem of recovering a block-sparse matrix from linear measurements.
In particular, denote $\textbf{B}=\begin{bmatrix}\mathbf{b}_{1}^{*}\\ \vdots\\ \mathbf{b}_{N}^{*}\end{bmatrix}=\begin{bmatrix}\mathbf{c}_{1}&\cdots&\mathbf{c}_{k}\end{bmatrix}\in\mathbb{C}^{N\times k}$, $\mathbf{\Psi}^{*}=\begin{bmatrix}\mathbf{q}_{1}^{T}\\ \vdots\\ \mathbf{q}_{N}^{T}\end{bmatrix}=\begin{bmatrix}\boldsymbol{\psi}_{1}&\cdots&\boldsymbol{\psi}_{N}\end{bmatrix}\in\mathbb{C}^{N\times N}$.
@@
$\textbf{y}_{i}=\textbf{F}_{\Omega}\begin{bmatrix}\textbf{b}_{1}^{*}\textbf{h}_{i}\textbf{z}^{T}\textbf{q}_{1}\\ \textbf{b}_{2}^{*}\textbf{h}_{i}\textbf{z}^{T}\textbf{q}_{2}\\ \vdots\\ \textbf{b}_{N}^{*}\textbf{h}_{i}\textbf{z}^{T}\textbf{q}_{N}\end{bmatrix}=\mathbf{F_{\Omega}}\underbrace{\begin{bmatrix}(\mathbf{\bar{q}}_{1}\bigotimes\mathbf{b}_{1})^{*}\\ \vdots\\ \mathbf{\bar{q}}_{N}\bigotimes\mathbf{b}_{N})^{*}\end{bmatrix}}_{\mathbf{\Phi}}\text{vec}(\textbf{h}_{i}\textbf{z}^{T}):=\mathbf{F}_{\Omega}\mathbf{\Phi}\text{vec}(\textbf{h}_{i}\textbf{z}^{T}),i=1\cdots C,$ (23)
where
$\boldsymbol{\Phi}\coloneq\begin{bmatrix}(\mathbf{\bar{q}}_{1}\bigotimes\mathbf{b}_{1})^{*}\\ \vdots\\ \mathbf{\bar{q}}_{N}\bigotimes\mathbf{b}_{N})^{*}\end{bmatrix}\in\mathbb{C}^{N\times kN}$ (24)
Defining
$\textbf{Y}=\begin{bmatrix}\textbf{y}_{1}|\textbf{y}_{2}|\cdots|\textbf{y}_{C}\end{bmatrix},$
$\textbf{H}=\begin{bmatrix}\textbf{h}_{1}|\textbf{h}_{2}|\cdots|\textbf{h}_{C}\end{bmatrix},$
$\textbf{X}=\begin{bmatrix}\textbf{X}_{1}|\textbf{X}_{2}|\cdots|\textbf{X}_{C}\end{bmatrix}=\begin{bmatrix}\text{vec}(\textbf{h}_{1}\textbf{z}^{T})|\text{vec}(\textbf{h}_{2}\textbf{z}^{T})|\cdots|\text{vec}(\textbf{h}_{C}\textbf{z}^{T})\end{bmatrix}=\textbf{z}\otimes\textbf{H}\in\mathbf{C}^{kN\times C},$
we can rewrite (23) in matrix form as follows:
@@
where $\textbf{Y}\in\mathbb{C}^{L\times C}$, $\mathbf{F}_{\Omega}\in\mathbb{C}^{L\times N}$, $\mathbf{\Phi}\in\mathbb{C}^{N\times kN}$, $\textbf{X}\in\mathbb{C}^{kN\times 1}$. We refer to $\mathbf{A}=\mathbf{F}_{\Omega}\mathbf{\Phi}$ as the forward matrix or measurement ensemble. Now the problem is to find the matrix $\mathbf{X}$ satisfying the linear constraint described by (25). To that end we will exploit the block sparsity of X.
@@ 3.2 Block sparsity of $\mathbf{X}$
Moreover, X is at most rank-k if $C$ is large. Methods that exploit the low-rankness include [Akçakaya_Basha_Goddu_Goepfert_Kissinger_Tarokh_Manning_Nezafat_2011, trzasko_manduca_2011, Holme_Rosenzweig_Ong_Wilke_Lustig_Uecker_2019] and also the the ENLIVE formulation (17) introduced in Section 2.2. Naturally, one could minimize over a linear combination $\|\mathbf{\tilde{X}}\|_{1,2}+\lambda\|\mathbf{\tilde{X}}\|_{*},\lambda>0$ in an attempt to simultaneously capture the low-rankness and sparse characteristics of $\mathbf{\tilde{X}}$. However, as demonstrated in [Oymak_Jalali_Fazel_Eldar_Hassibi_2015], this approach is no more effective than using a single norm in terms of the number of measurements required. Theoretically, only recovering for a single coil case by minimizing $\|\cdot\|_{*}$ requires $L=O(N+k)$. Also $l_{1,2}$ minimization is more computationally efficient than the nuclear norm minimization problem. Hence, we choose to only minimize with respect to $\|\cdot\|_{1,2}$, which (as we will prove) may already yield a solution with sparse structure and small nuclear norm.
@@ 3.3 Optimization problem
which is the sum of the Frobenius norm of the $k$ by $C$ blocks of $\mathbf{X}$.
Since it is well known that the $\ell_{1,2}$-norm promotes the block sparse structure in $\mathbf{X}$ [eldar_kuppinger_bolcskei_2010, eldar_mishali_2009, flinth_2017], we solve the following convex program
@@
If the measurement are noisy, i.e., $\textbf{Y}=\textbf{F}_{\Omega}\mathbf{\Phi}\textbf{X}+\mathbf{N}$ with $\|\mathbf{N}\|_{F}\leq\sqrt{C}\sigma$, we solve the following convex program
$\min\|\mathbf{X}\|_{1,2}\quad\text{subject to}\|\textbf{Y}-\textbf{F}_{\Omega}\mathbf{\Phi}\textbf{X}\|_{F}\leq\sqrt{C}\sigma.$ (28)
@@
Once we get a solution $\hat{\textbf{X}}$ to the $\ell_{1,2}$ minimization problem (27)(28), we calculate the average of the best rank-one estimations of its columns to retrieve the common underlying signal. Recall the columns of $\hat{\textbf{X}}=\begin{bmatrix}\hat{\textbf{x}}_{1}&\cdots\hat{\textbf{x}}_{C}\end{bmatrix}$ and the truth $\textbf{X}^{0}=\begin{bmatrix}\textbf{X}^{0}_{1}&\cdots\textbf{X}^{0}_{C}\end{bmatrix}=\begin{bmatrix}\text{vec}(\textbf{h}_{1}\textbf{z}^{T})|\text{vec}(\textbf{h}_{2}\textbf{z}^{T})|\cdots|\text{vec}(\textbf{h}_{C}\textbf{z}^{T})\end{bmatrix}$. We use the following Lemma from [ling_strohmer_2015] to compute an average of the estimates from $\hat{\mathbf{X}}$ and to ensure that it is close to z.
@@
For any solution $\textbf{X}=\begin{bmatrix}\textbf{x}_{1}|\cdots|\textbf{x}_{C}\end{bmatrix}$ and the lifted matrix $\textbf{X}_{0}=\begin{bmatrix}\text{vec}(\textbf{h}_{1}\textbf{z}_{0}^{T})|\text{vec}(\textbf{h}_{2}\textbf{z}_{0}^{T})|\cdots|\text{vec}(\textbf{h}_{C}\textbf{z}_{0}^{T}))\end{bmatrix}$, let $\hat{\sigma}\hat{\textbf{u}}\hat{\textbf{v}}^{T}$ be the best rank-one Frobenius norm approximation of $\sum\limits_{i}\textbf{x}_{i}/C$ reshaped into a matrix $\hat{\mathbf{X}}$ $\in\mathbb{C}^{k\times N}$ in the column-wise order. If $\|\textbf{X}-\textbf{X}_{0}\|_{F}=\sqrt{C}\epsilon$, then there exists a scalar $\alpha$ and a constant $C_{0}$ such that,
$\|\textbf{z}_{0}-\alpha^{-1}\hat{\textbf{v}}\|\leq C_{0}\min(\epsilon/\|\textbf{z}_{0}\|,\|\textbf{z}_{0}\|).$
The Lemma is a direct consequence of combining Corollary 3.3 from [ling_strohmer_2015] and using norm equivalency. Also notice if for any $\alpha_{0}\neq 0$, $\alpha_{0}\textbf{u}_{0}$ and $1/\alpha_{0}\textbf{v}_{0}$ is also a pair of solutions. However, the scalar ambiguity is not important for many applications, including pMRI. The reconstructed signal is retrieved as $\hat{\textbf{x}}=\boldsymbol{\Psi}^{*}\hat{\textbf{v}}$
@@
A comprehensive review of the algorithms to solve problems of the form (29) is not a focus of this paper. We refer to a few common approaches such as the forward-backward splitting (FBS) [Combettes_Patrick_Wajs_Valerie_2005], ADMM [wang_yang_yin_zhang_2008] and FISTA [Beck_Teboulle_2009] which have found numerous applications in the context of pMRI reconstruction. In the simulations part of this paper, we use a FISTA-type iterative method to solve the unconstrained problem. Detailed iterative updates are provided in the Appendix 7.2.
@@ 4 Main result
Given any integer $k$ and $N$, we define $T_{j}\subset[kN]$ by $T_{j}=\{(j-1)k+1,\cdots,jk\},j\in[N]$. Given an index set $S=\{s_{1},s_{2},\cdots\}\subset[N]$, define $S_{j}=T_{s_{j}}$ for $j\in|S|$ and $T:=\cup_{j\in|S|}S_{j}$. With slight abuse of notation, we will not distinguish $\mathcal{P}_{S}$, $\mathcal{P}_{T}$ and $\mathcal{P}_{T\times[C]}$ when it is clear from the context. For example, for $\textbf{x}\in C^{N}$, $\textbf{x}\in C^{kN}$ and $\textbf{X}\in C^{kN\times C}$, $\mathcal{P}_{S}$ means the projection onto $S$, $T$ and $T\times[C]$ respectively. We will also sometimes write the projections $\mathcal{P}_{S}$ as subscripts, for example, $\textbf{x}_{S}$ or $\textbf{X}_{S}$, when it is convenient.
@@
It will come in handy in the proofs to decompose a linear map to any given subset of its columns. To that end, for a matrix $\textbf{A}\in\mathbb{C}^{L\times kN}$ and a given set $S$, we define the projections onto its blocks of columns induced by $S$ as